OpenStudy (anonymous):

I Will Medal! What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (igreen):

\(\sf 5(8 + \color{red}7) = 5 \cdot x + 5 \cdot \color{red}7\)

OpenStudy (igreen):

When distributing, we multiply the number outside the parenthesis to all the numbers inside the parenthesis. \(\sf \color{lime}5(8 + \color{red}7) = \color{lime}5 \cdot x + \color{lime}5 \cdot \color{red}7\)

OpenStudy (igreen):

So what's 'x' going to be?

OpenStudy (igreen):

\(\sf \color{lime}5(\color{blue}8 + \color{red}7) = \color{lime}5 \cdot \color{blue}x + \color{lime}5 \cdot \color{red}7\)

OpenStudy (acxbox22):

remember 5(8+7)=5(8)+5(7) ~ ~ (x)